### v2

# seq-JEPA: Autoregressive Predictive Learning of Invariant-Equivariant World Models

Abstract
Current self-supervised algorithms commonly rely on transformations such as data augmentation and masking to learn visual representations. This is achieved by enforcing invariance or equivariance with respect to these transformations after encoding two views of an image. This dominant two-view paradigm often limits the flexibility of learned representations for downstream adaptation by creating performance trade-offs between high-level invariance-demanding tasks such as image classification and more fine-grained equivariance-related tasks. In this work, we proposes seq-JEPA, a world modeling framework that introduces architectural inductive biases into joint-embedding predictive architectures to resolve this trade-off. Without relying on dual equivariance predictors or loss terms, seq-JEPA simultaneously learns two architecturally segregated representations: one equivariant to specified transformations and another invariant to them. To do so, our model processes short sequences of different views (observations) of inputs. Each encoded view is concatenated with an embedding of the relative transformation (action) that produces the next observation in the sequence. These view-action pairs are passed through a transformer encoder that outputs an aggregate representation. A predictor head then conditions this aggregate representation on the upcoming action to predict the representation of the next observation. Empirically, seq-JEPA demonstrates strong performance on both equivariant and invariant benchmarks without sacrificing one for the other. Furthermore, it excels at tasks that inherently require aggregating a sequence of observations, such as path integration across actions and predictive learning across eye movements.
1 Introduction
Self-supervised learning (SSL) in latent space has made significant progress in visual representation learning, closing the gap with supervised methods across many tasks. Most SSL methods rely on comparing two transformed views of an image and enforcing invariance to the transformations (Misra and van der Maaten, 2020; Chen et al., 2020; He et al., 2020; Dwibedi et al., 2021; HaoChen et al., 2021; Yeh et al., 2022; Caron et al., 2020, 2021; Ermolov et al., 2021; Assran et al., 2022; Zbontar et al., 2021; Bardes et al., 2022). Another group of methods employ techniques to preserve transformation-specific information, thereby learning equivariant representations (Lee et al., 2021; Xiao et al., 2021; Park et al., 2022; Dangovski et al., 2022; Gupta et al., 2023; Garrido et al., 2023, 2024; Gupta et al., 2023, 2024; Yerxa et al., 2024).
Equivariance is a crucial representational property for downstream tasks that require fine-grained distinctions. For example, given representations that are invariant to color, it is not possible to distinguish between certain species of flowers or birds (Lee et al., 2021; Xiao et al., 2021). Moreover, recent work has shown that equivariant representations are better aligned with neural responses in primate visual cortex and could be important for building more accurate models thereof (Yerxa et al., 2024). While some equivariant SSL approaches have reported minor gains on tasks typically associated with invariance (e.g., classification) (Devillers and Lefort, 2022; Park et al., 2022; Gupta et al., 2023), a growing body of work highlights a fundamental trade-off between learning invariance and equivariance, i.e., models that capture equivariance-related style latents do not fare well in classification and vice versa (Garrido et al., 2023, 2024; Gupta et al., 2024; Yerxa et al., 2024; Rusak et al., 2025). This trade-off has recently received theoretical support (Wang et al., 2024), underscoring the need for new architectural and objective designs that can reconcile these competing goals.
In contrast to the two-view paradigm in SSL, humans and other animals rely on a sequence of actions and consequent observations (views) for developing appropriate visual representation during novel object learning  (Harman et al., 1999; Vuilleumier et al., 2002). For example, they recognize a 3-D object by changing their viewpoint and examining different sides of the object (Tarr et al., 1998). Inspired by this, we introduce seq-JEPA, a self-supervised world modeling framework that combines joint-embedding predictive architectures (LeCun, 2022) with inductive biases for sequential processing. seq-JEPA simultaneously learns two architecturally distinct representations: one that is equivariant to a specified set of transformations, and another that is suited for invariance-demanding tasks, such as image classification.
Specifically, our framework (Figure 1) processes a short sequence of transformed views (observations) of an image. Each view is encoded and concatenated with an embedding corresponding to the relative transformation (action) that produces the next observation in the sequence. These view-action pairs are passed through a transformer encoder, a form of learned working memory, that outputs an aggregate representation of them. A predictor head then conditions this aggregate representation on the upcoming action to predict the representation of the next observation.
Our results demonstrate that individual encoded views in seq-JEPA become transformation/action-equivariant. Through ablations, we show that action conditioning plays a key role in promoting equivariant representation learning in the encoder network. In contrast, the aggregate representation of views, produced at the output of the transformer, becomes largely transformation/action-invariant. This emergent architectural disentanglement of invariant and equivariant representations is central to seq-JEPA’s competitive performance compared to invariant and equivariant SSL methods on both categories of tasks (Figure 3). Unlike most prior equivariant SSL methods (Lee et al., 2021; Park et al., 2022; Dangovski et al., 2022; Gupta et al., 2023; Garrido et al., 2023; Gupta et al., 2024; Yerxa et al., 2024), our model does not rely on explicitly crafted loss terms or objectives to achieve equivariance, nor is it instructed to learn the decomposition of two representation types. Instead, the dual representation structure arises naturally from the model architecture and action-conditioned predictive learning.
Beyond resolving the invariance-equivariance trade-off, seq-JEPA further benefits from moving away from the two-view paradigm in SSL. We show that it performs well on tasks requiring integration over sequences of observations. In one scenario, inspired by embodied vision in primates, our model learns image representations without augmentations or masking, solely by predicting across simulated eye movements (saccades). In another setting, it performs path integration over sequences of actions—such as eye movements or 3D object rotations in 3DIEBench (Garrido et al., 2023). Our key contributions are as follows:
  • We introduce seq-JEPA, a self-supervised world model that learns architecturally distinct equivariant and invariant representations through sequential prediction over action-observation pairs, without requiring explicit equivariance losses or dual predictors.
  • We empirically validate that seq-JEPA matches or outperforms existing invariant and equivariant SSL methods across tasks requiring either representational property.
  • We demonstrate that seq-JEPA naturally supports tasks that involve sequential integration of observations, such as predictive learning across saccades and path integration over action sequences.
2 Method
2.1 Invariant and equivariant representations
Before presenting our architecture and training procedure, we briefly define invariance and equivariance in the context of SSL (Dangovski et al., 2022; Devillers and Lefort, 2022). Let $\mathcal{T}$ denote a distribution over transformations, parameterized by a vector $t$. These transformations—such as augmentations or masking—can be used to generate multiple views from a single image $x$. Let $x_{1}$ and $x_{2}$ be two such views, produced by applying transformations $t_{1}$ and $t_{2}$ sampled from $\mathcal{T}$. Additionally, let $a$ denote the relative transformation that maps $x_{1}$ to $x_{2}$. Additionally, we denote $a$ as a transformation that transforms $x_{1}$ to $x_{2}$. We distinguish between $t$ (an individual transformation) and $a$ (an action), where the latter reflects the change from one view to another. Let $f$ be an encoder that maps inputs to a latent space. We say that $f$ is equivariant to $t$ if:
$\forall{t\in\mathcal{T}},~{}\exists{,u_{t}};\mathrm{s.t.}\qquad f(t(x))=u_{t}(f(x)),$ (1)
where $u_{t}$ is a transformation in latent space corresponding to $t$. Equivariance can similarly be defined in terms of relative transformations (actions):
$\forall{a\in\mathcal{T}},~{}\exists{,u_{a}};\mathrm{s.t.}\qquad f(x_{2})=u_{a}(f(x_{1})).$ (2)
As a special case, if $u_{t}$ and $u_{a}$ are identity functions, then $f$ is invariant to the transformation: $f(t(x))=f(x)$ or $f(x_{2})=f(x_{1})$.
2.2 Architecture
Figure 1 presents the overall architecture of seq-JEPA. Let $\{x_{i}\}_{i=1}^{M+1}$ be a sequence of views generated from a sample $x$ via transformations $\{t_{i}\}_{i=1}^{M+1}$. The relative transformations (actions) $\{a_{i}\}_{i=1}^{M}$ are defined as $a_{i}\triangleq\Delta t_{i,i+1}$, i.e., the transformation mapping $x_{i}$ to $x_{i+1}$. In our default setting, we use a learnable linear projector to encode these actions.
A backbone encoder, $f$ encodes the first $M$ views, producing representations $\{z_{i}\}_{i=1}^{M}$. Except for $z_{M}$, each $z_{i}$ is concatenated with its corresponding action embedding and passed to a transformer encoder $g$ (no MLP projector is used after the encoder), which aggregates the sequence of action-observation pairs. The transformer uses a learnable [AGG] token (analogous to the [CLS] token in ViTs (Dosovitskiy et al., 2020)) to generate the aggregate representation:
$z_{AGG}=g((z_{1},a_{1}),(z_{2},a_{2})...,(z_{M-1},a_{M-1}),z_{M})$ (3)
This aggregate representation $z_{AGG}$ is then concatenated with the final action embedding $a_{M}$ (corresponding to the transformation from $x_{M}$ to $x_{M+1}$), and passed to an MLP predictor $h$ to predict the representation of $x_{M+1}$:
$\hat{z}_{M+1}=h(z_{AGG},a_{M}).$ (4)
The ground truth $z_{M+1}$ is computed using a target encoder—an exponential moving average (EMA) of $f$. The target representation is passed through a stop-gradient operator (sg) to avoid representational collapse. The training objective is to maximize the cosine similarity between $\hat{z}_{M+1}$ and $z_{M+1}$:
$\mathcal{L}_{\mathrm{seq-JEPA}}=1-\frac{\hat{z}_{M+1}}{\left\lVert{\hat{z}_{M+1}}\right\rVert_{2}}\cdot\frac{\texttt{sg}(z_{M+1})}{\left\lVert{\texttt{sg}(z_{M+1})}\right\rVert_{2}}.$ (5)
No additional loss terms or equivariance-specific predictors are used during training.
Figure 1: seq-JEPA is a world modeling paradigm that leverages a sequence of action-observation pairs to learn architecturally distinct invariant and equivariant representations.
Figure 2: a. Transformations and observations used for training b. In predictive learning across saccades, image saliencies and inhibition of return help create more informative and less redundant patch sequences.
2.3 Action and observation sets
To evaluate generalization across transformation types, we define three sets of action-observation pairs (Figure 2). See Appendix A.1 for details of each setup.
3D Invariant Equivariant Benchmark (3DIEBench). We first evaluate on the 3DIEBench dataset (Garrido et al., 2023), designed for equivariance analysis. It includes 3D object renderings with variations in rotation, floor hue, and lighting. We primarily study equivariance to $SO(3)$ rotations and secondarily to appearance factors of floor and light hue.
Hand-Crafted Augmentations. In this setting, we use transformed views generated via common SSL augmentations (e.g., crop, color jitter, blur), where actions correspond to relative augmentation parameters. We use CIFAR100 and Tiny ImageNet, and follow EquiMod’s augmentation protocol (Devillers and Lefort, 2022).
Predictive Learning Across Saccades (PLS). Going beyond conventional transformations such as augmentations or 3D rotations, we show that seq-JEPA can learn visual representations from a sequence of partial observations thanks to architectural inductive biases—without relying on any hand-crafted augmentations. Our PLS has a similar flavor to I-JEPA (Assran et al., 2023) but does not require engineered masking strategies. In PLS, we train seq-JEPA on sequences of patches extracted from high-resolution images. For instance, with STL-10, we use $32\times 32$ patches to form the observation sequence. In this setting, actions correspond to the relative positions between patch centers, simulating saccadic eye movements and inducing 2-D positional equivariance to representations. To select fixation points, we adopt two biologically inspired techniques that increase informativeness, reduce redundancy, and improve the downstream utility of the aggregate representation (Figure 2):
  • Saliency-Based Fixation Sampling. Using DeepGaze IIE (Linardos et al., 2021), we extract saliency maps for each image and use them to probabilistically sample fixation points (Itti et al., 1998; Li, 2002; Zhaoping, 2014). The maps are pre-computed and introduce no training overhead.
  • Inhibition of Return (IoR). To reduce spatial overlap between patches and emulate natural exploration (Posner et al., 1985), we implement IoR by zeroing out the sampling probability of areas surrounding previously sampled fixations.
3 Experimental Setup
3.1 Compared methods and baselines
We compare seq-JEPA against both invariant and equivariant SSL baselines. Invariant methods include SimCLR (Chen et al., 2020), BYOL (Grill et al., 2020), and VICReg (Bardes et al., 2022). Equivariant methods include SEN (Park et al., 2022), EquiMod (Devillers and Lefort, 2022), SIE (Garrido et al., 2023), and ContextSSL (Gupta et al., 2024). For all baselines, architectural details are given in Appendix A.3. We also evaluate two hybrid baselines based on our architecture:
  • Conditional BYOL. A two-view version of seq-JEPA with no sequence aggregator, where BYOL’s predictor is conditioned on the relative transformation between target and online views. This encourages representations to encode transformation information.
  • Conv-JEPA. A baseline for the saccades setting. It uses the same sequence of saliency-sampled patches as seq-JEPA and predicts the final patch’s representation from each earlier patch individually. These losses are summed across the pairs before backpropagation.
3.2 Training protocol
All models use ResNet-18 (He et al., 2016) as the backbone encoder. For action conditioning, we use a learnable linear projection to learn action embeddings (default action embedding is 128-d). The sequence aggregator in seq-JEPA is a lightweight transformer encoder (Vaswani et al., 2017) with three layers and four attention heads. The predictor is a 2-layer MLP with 1024 hidden units and ReLU activation. In order to control for and eliminate any performance gain resulting from using a transformer encoder in seq-JEPA instead of an MLP projection head, we train baselines that typically use MLP projectors in two variants: (1) with original MLP projector; and (2) with the MLP replaced by a transformer encoder and a sequence length of one. We report the better-performing variant to ensure a fair comparison. All models are trained from scratch with a batch size of 512. We use 1000 epochs for 3DIEBench and 2000 epochs for other datasets to obtain asymptotic performance. We use AdamW for models with transformer projectors (including seq-JEPA) due to its stability and improved regularization in transformer training. For ConvNet-only models with MLP heads, we use the Adam optimizer. Full hyperparameters are detailed in Appendix A.
3.3 Evaluation metrics and protocol
To assess equivariance, we follow the protocol of Garrido et al. (2023) and train a regressor on frozen encoder representations to predict the relative transformation (action) between two views. We report the $R^{2}$ score on the test set. As a proxy measure of invariance, we use top-1 classification accuracy of a linear probe on top of frozen representations. For all baselines, probes are trained on encoder outputs. For seq-JEPA, we measure accuracies on top of the aggregate representation ($z_{AGG}$ in Figure 1) and report the number of observation views used during training and inference. For completeness, we also report classification performance on top of encoder representation ($z_{i}$ in Figure 1) for seq-JEPA models in Appendix B.4. Training details of evaluation heads are given in Appendix A.2.
4 Results
4.1 Quantitative evaluation on 3DIEBench
We use the 3DIEBench benchmark to quantitatively compare invariant and equivariant representations in seq-JEPA with baseline methods. This benchmark allows us to measure equivariance through decoding 3D object rotations while enabling invariance measurement through object classification. Table 3 provides a summary of our evaluation on the 3DIEBench where equivariant methods have been conditioned on rotation. In addition to the relative rotation between two views, in the last column we provide the $R^{2}$ score for predicting individual transformation parameters from representations of a single view. For seq-JEPA, we trained models with varying training sequence lengths (denoted by $M_{tr}$ in the table) and measure the linear classification performance on top of aggregate representations with different inference lengths (denoted by $M_{val}$).
Among invariant methods, BYOL achieves the highest classification accuracy, yet does not offer a high level of equivariance. Among equivariant baselines, SIE and ContextSSL yield strong rotation prediction performance due to their specialized equivariance predictors and loss functions, but underperform in classification. EquiMod and SEN offer better classification performance, yet compromise equivariance. In contrast, seq-JEPA achieves strong performance in both, matching the best rotation $R^{2}$ scores while exceeding baselines in classification, with gains increasing with inference sequence length $M_{val}$. Ablating action conditioning leads to a sharp drop in equivariance but retains classification accuracy, confirming our hypothesis that action-conditioned sequential aggregation enables a segregated invariant-equivariant representation structure. For additional results on 3DIEBench including models with varying training and inference sequence lengths and models conditioned on both rotation and color, see Appendices B.1 to B.3.
Figure 3: Top-1 linear classification (invariance) vs. rotation prediction (equivariance) performance on 3DIEBench. seq-JEPA learns good representations for both tasks. Subscripts indicate training and inference sequence lengths.
Table 1: Evaluation on 3DIEBench for linear probe classification (invariance) and rotation prediction (equivariance). Equivariant models and seq-JEPA are conditioned on rotation. For seq-JEPA, training and inference sequence lengths are denoted by $M_{tr}$ and $M_{val}$.
Method Top-1 Acc. (%) Rel. Rot. ($R^{2}$) Indiv. Rot. ($R^{2}$)
Invariant
BYOL 82.90 0.12 0.25
SimCLR 81.13 0.35 0.54
VICReg 80.48 0.20 0.36
Equivariant
SEN 83.43 0.35 0.57
EquiMod 84.29 0.32 0.55
SIE 77.49 0.58 0.62
Conditional BYOL 82.61 0.31 0.47
ContextSSL ($c=126$) 80.40 0.74 0.78
Invariant-Equivariant
seq-JEPA (1, 1) 84.08 0.65 0.69
seq-JEPA (1, 3) 85.31 0.65 0.69
seq-JEPA (3, 3) 86.14 0.71 0.74
seq-JEPA (3, 5) 87.41 0.71 0.74
seq-JEPA (no act cond) 86.05 0.29 0.37
4.2 Qualitative evaluation on 3DIEBench
To visualize equivariance in representational space, we retrieve the three nearest representations of a query image from the validation set of 3DIEBench (Figure 5). While all models retrieve the correct object category, only seq-JEPA and SIE consistently preserve rotation across all retrieved views, consistent with their high $R^{2}$ scores. Next, we projected encoder and aggregate representations using 2D UMAP (Figure 5). The left panel shows encoder representations colored by class label, while the middle panel displays the same encoder representations colored by rotation angle. The smooth color gradation across the map within each class cluster in the middle panel suggests that the encoder captures rotation angle as a continuous factor, implying equivariance to rotation (e.g., the red class in the bottom-right corner of the right panel and the corresponding part in the middle panel). The right panel shows aggregate representations colored by class label. Comparing the class-colored plots (left and right panels), we observe that both encoder and aggregate representations contain class information. However, when we aggregate multiple views of a sample, some of the intra-class variability (resulting from transformations such as rotation) is eliminated, causing each class’ representational cluster to become more homogeneous. This aggregation procedure likely reduces variation due to rotation and makes the representations more invariant, resulting in decreased intra-class spread and increased inter-class distance. We create a similar UMAP visualization for seq-JEPA with ablated rotation conditioning in Appendix B.5 to highlight the role of action conditioning in achieving equivariance to rotation.
Figure 4: Retrieval of nearest representations; given a query image, we extract the three nearest encoder representations in the validation set of 3DIEBench. The retrieved views of models with the highest quantitative rotation equivariance performance maintain the rotation of the query image across all retrieved views.
Figure 5: 2-D UMAP projections of seq-JEPA’s encoder and aggregate representations trained on 3DIEBench and conditioned on rotation with $M_{tr}=3$ and $M_{val}=5$. Encoder representations for each view observation, color-coded by class (left) and rotation angle (middle). Aggregate representation for $M_{val}=5$, color-coded by class (right).
4.3 Evaluation with Hand-Crafted Augmentations
We assess invariance and equivariance under hand-crafted augmentations by training on CIFAR100 and Tiny ImageNet (Table 2). Models with action conditioning are trained conditioned on crop, color jitter, blur, or all three (indicated in the first column of the table). seq-JEPA consistently achieves higher equivariance than both invariant and equivariant baselines across all transformations. Notably, except for the model trained on CIFAR-100 and conditioned on blur, the best equivariance performance for a given augmentation is achieved when the model is specialized and conditioned only on that augmentation. Furthermore, ablating actions (last row in the table) causes seq-JEPA to lose its equivariance across transformations compared to action-conditioned models. Overall, our model outperforms both invariant and equivariant families in terms of equivariance, while being competitive in terms of classification performance. For additional results with varying training and inference sequence lengths, see Appendix B.3.
Table 2: Evaluation with hand-crafted augmentations on CIFAR100 and Tiny ImageNet; equivariance is measured by predicting relative transformation parameters associated with crop, color jitter, or blur augmentations. For all seq-JEPA models, $M_{val}=5$.
CIFAR100 Tiny ImageNet
Conditioning Method Classification (top-1) Crop ($R^{2}$) Jitter ($R^{2}$) Blur ($R^{2}$) Classification (top-1) Crop ($R^{2}$) Jitter ($R^{2}$) Blur ($R^{2}$)
Invariant
- SimCLR 61.72 0.56 0.18 0.04 34.29 0.30 0.08 0.13
BYOL 62.17 0.39 0.05 -0.01 35.70 0.17 0.01 0.16
VICReg 61.35 0.49 0.11 -0.06 35.29 0.31 0.05 0.19
Equivariant
Crop+Jitter+Blur SEN 61.94 0.65 0.51 0.85 36.01 0.24 0.48 0.87
EquiMod 61.80 0.59 0.49 0.74 36.75 0.38 0.46 0.86
SIE 58.81 0.34 0.26 0.53 31.37 0.26 0.56 0.88
Conditional BYOL 60.63 0.56 0.46 0.73 35.49 0.34 0.54 0.87
Crop SEN 61.56 0.66 0.15 0.10 35.95 0.24 0.11 0.49
EquiMod 61.77 0.62 0.15 0.01 36.83 0.26 0.12 0.30
SIE 57.55 0.69 0.11 0.25 32.38 0.34 0.05 0.14
Conditional BYOL 60.17 0.55 0.10 -0.01 37.07 0.22 0.01 0.24
Color jitter SEN 61.78 0.50 0.52 0.02 36.59 0.26 0.50 0.21
EquiMod 61.53 0.44 0.50 -0.02 37.18 0.20 0.52 0.29
SIE 59.29 0.48 0.59 0.06 34.37 0.39 0.62 0.35
Conditional BYOL 61.30 0.36 0.52 0.02 37.93 0.24 0.62 0.30
Blur SEN 61.47 0.43 0.15 0.84 34.62 0.16 0.08 0.79
EquiMod 62.72 0.41 0.15 0.74 36.12 0.24 0.10 0.91
SIE 57.66 0.40 0.07 0.71 31.00 0.26 0.05 0.85
Conditional BYOL 60.74 0.36 0.11 0.69 36.17 0.19 0.02 0.85
Invariant-Equivariant
Crop+Jitter+Blur seq-JEPA ($M_{tr}=1$) 52.90 0.77 0.52 0.23 37.10 0.64 0.49 0.89
Crop+Jitter+Blur seq-JEPA ($M_{tr}=2$) 60.17 0.78 0.64 0.88 35.56 0.69 0.42 0.93
Crop+Jitter+Blur seq-JEPA ($M_{tr}=3$) 58.33 0.79 0.63 0.92 34.85 0.67 0.64 0.96
Crop seq-JEPA ($M_{tr}=2$) 59.32 0.78 0.01 0.10 35.74 0.70 0.12 0.46
Color Jitter seq-JEPA ($M_{tr}=3$) 58.62 0.68 0.68 0.29 35.21 0.60 0.66 0.62
Blur seq-JEPA ($M_{tr}=3$) 56.82 0.71 0.15 0.74 35.79 0.58 0.22 0.97
- seq-JEPA ($M_{tr}=2$) 58.37 0.64 0.14 0.16 35.97 0.52 0.18 0.47
4.4 Predictive Learning across Saccades and Path Integration
In our third action-observation setting, we consider predictive learning across simulated eye movements to exhibit seq-JEPA’s ability in leveraging a sequence of partial observations to learn visual representations in a predictive manner. Table 6 shows that seq-JEPA reaches 83.44% top-1 accuracy on STL-10, comparable to SimCLR (85.23%) trained with full-resolution images and strong augmentations. This gap narrows further when increasing the inference sequence length from $M_{val}=4$ to $M_{val}=6$. Ablating action conditioning causes the accuracy on top of the aggregate representations to drop sharply, indicating that 2-D positional awareness is essential to forming semantic representations across simulated eye movements. Compared to Conv-JEPA, which accumulates prediction losses pairwise, seq-JEPA performs better in classification, highlighting the importance of sequence aggregation when dealing with partial observations in SSL. Further ablations show that saliency-driven sampling and IoR are critical for forming informative, non-overlapping patch sequences and subsequently a high-quality aggregate representation. Interestingly, while random uniform patch sampling negatively impacts classification accuracy due to lower semantic content, it results in the highest positional equivariance as the model samples patches and corresponding saccade actions from a more diverse set of positions across the entire image, not just the salient regions. UMAP projections for PLS with and without action conditioning in Appendix B.5 further underscore the role of action conditioning in enabling positional equivariance.
Path Integration. In the context of eye movement-driven observations and in general sequential observations, another ability that naturally arises is path integration (McNaughton et al., 2006)—predicting the cumulative transformation/action from a sequence of actions. We evaluate this task in both eye movements in PLS (visual path integration) and object rotations in 3DIEBench (angular path integration). As shown in Figure 6, seq-JEPA demonstrates strong performance in both settings, with performance degrading gracefully as sequence length increases. Ablating action conditioning causes path integration to fail, whereas ablating the visual stream has only a minor impact—highlighting that action information is the dominant signal for this task. Full details of the path integration setup are provided in Appendix B.6.
Table 3: Evaluation of predictive learning across saccades on STL-10. Equivariance is measured with respect to fixation coordinates. Unless stated otherwise, $M_{tr}=M_{val}=4$.
Conditioning Top-1 Acc. (%) Position ($R^{2}$)
Invariant
- SimCLR (augmentations) 85.23 -0.06
Equivariant
position Conv-JEPA ($M_{tr,val}=4$) 80.04 0.80
Invariant-Equivariant
- seq-JEPA 70.45 0.38
position seq-JEPA 83.44 0.80
position seq-JEPA ($M_{val}=6$) 84.12 0.80
position seq-JEPA (w/o saliency & IoR) 79.85 0.88
position seq-JEPA (w/o IoR) 77.97 0.85
Figure 6: a) Visual path integration across eye movements, b) Angular path integration across object rotations (results over three random seeds)
4.5 Ablations: Action Conditioning Is Key
To better understand the mechanisms underlying seq-JEPA’s invariant-equivariant representation learning and role of action conditioning, we perform a set of ablation experiments on 3DIEBench. Specifically, we study: (i) the role of action conditioning in the transformer and predictor; (ii) the impact of learning action embeddings and their dimensionality. Table 7 summarizes our ablation results. Removing action conditioning entirely causes a significant drop in equivariance ($R^{2}$ from 0.71 to 0.29), although classification accuracy remains high thanks to sequence aggregation and segregated invariance-equivariance in our model. Conditioning only the transformer or only the predictor leads to intermediate results, with predictor conditioning proving more critical for equivariance. We also vary the dimensionality of the learnable action embeddings: performance saturates around the default size of 128, with smaller sizes (e.g., 16 or 64) already sufficient to capture the rotation structure. Non-learnable embeddings significantly reduce equivariance, underscoring the importance of learned representations in the action pathway.
Table 4: Ablation results for action conditioning (3DIEBench). All models use $M_{tr}=3$, $M_{val}=5$ (results over three random seeds).
Variant Top-1 Acc. (%) Rotation ($R^{2}$)
Act. conditioning
None 87.36 $\pm$ 0.7 0.29 $\pm$ 0.04
No predictor cond. 87.17 $\pm$ 0.3 0.37 $\pm$ 0.06
No transformer cond. 86.33 $\pm$ 0.1 0.53 $\pm$ 0.05
Act. embedding dim.
Non-learnable 85.68 $\pm$ 0.2 0.34 $\pm$ 0.04
$a_{\text{dim}}=16$ 86.29 $\pm$ 0.4 0.70 $\pm$ 0.01
$a_{\text{dim}}=64$ 87.11 $\pm$ 0.2 0.70 $\pm$ 0.02
$a_{\text{dim}}=128$ (default) 87.41 $\pm$ 0.5 0.71 $\pm$ 0.02
$a_{\text{dim}}=256$ 87.26 $\pm$ 0.6 0.72 $\pm$ 0.00
Figure 7: Effect of training and inference sequence length on seq-JEPA’s performance; left:: Equivariant performance ($R^{2}$) versus training sequence length; middle: Classification accuracy versus training sequence length; right: Classification accuracy versus inference sequence length.
4.6 Scaling Properties: Role of Training and Inference Sequence Lengths
We study the effect of both training and inference sequence lengths on performance across tasks, i.e., scalability of seq-JEPA in terms of context length (Figure 7). We draw inspiration from recent findings in large-scale foundation models, where increased training and inference context—whether in text (Brown et al., 2020; Touvron et al., 2023), vision (Zellers et al., 2021; Chen et al., 2021), or video (Bain et al., 2021; Arnab et al., 2021)—consistently leads to stronger representations. We observe that equivariance generally improves with longer training sequences (left panel), likely due to increased exposure to action-conditioned transitions. Classification performance on 3DIEBench and STL-10 (middle panel) similarly benefits. In contrast, on CIFAR100 and Tiny ImageNet with synthetic augmentations, longer training sequences slightly decrease classification performance. We hypothesize that leveraging and aggregating a sequence of action-observation pairs, i.e. seq-JEPA’s architectural inductive bias, is most effective in settings where the downstream task benefits from sequential observations. In the case of object rotations in 3DIEBench, seeing an object from multiple angles is indeed beneficial in recognizing the object’s category, which explains the improved classification accuracy with increased training sequence length. Similarly, in the case of predictive learning across saccades, each eye movement and its subsequent glance provides additional information that can be leveraged for learning a richer aggregate representation.
At inference time, all datasets benefit from longer context lengths ($M_{val}$), confirming that richer aggregate representations yield stronger performance (right plot). This scalability via sequence lengths opens avenues for efficient representation learning with small foveated patches in lieu of full-frame inputs, mirroring how foundation models scale with input tokens at test time. Together with our transfer learning results on ImageNet-1k (Appendix B.7), these findings suggest that seq-JEPA’s architectural inductive bias enables graceful scaling via longer sequence lengths.
5 Related Work
Non-Generative World Models and Joint-Embedding Predictive Architectures. Non-generative world models predict the effect of transformations or actions directly in latent space, avoiding reconstruction in pixel space. This includes contrastive SSL methods that model transformed views from context representations (van den Oord et al., 2019; Gupta et al., 2024), as well as approaches in model-based reinforcement learning (RL) to improve sample efficiency, generate intrinsic rewards, or capture environment transitions (Schwarzer et al., 2021; Khetarpal et al., 2025; Ni et al., 2024; Tang et al., 2023; Guo et al., 2022). Joint-embedding predictive architectures (LeCun, 2022) form a structured subclass of non-generative world models. Rather than aligning or contrasting representations directly, they introduce an asymmetric predictor conditioned on transformation parameters to infer the outcome of an action applied to a latent view. Examples include I-JEPA(Assran et al., 2023), which predicts masked regions from positional cues, and IWM (Garrido et al., 2024), which conditions on augmentation parameters. JEPAs have been recently extended to physical reasoning (Garrido et al., 2025) and offline planning (Sobal et al., 2025), illustrating the framework’s versatility in structured predictive representation learning.
Equivariant SSL. Equivariant SSL methods aim to retain transformation-specific information in the latent space, typically by augmenting invariant objectives with an additional equivariance term. Some approaches directly predict transformation parameters (Lee et al., 2021; Scherr et al., 2022; Gidaris et al., 2018; Gupta et al., 2024; Dangovski et al., 2022). Methods such as EquiMod (Devillers and Lefort, 2022) and SIE (Garrido et al., 2023) predict the effect of a transformation in latent space via a predictor in addition to their invariant objective. SEN (Park et al., 2022) similarly predicts transformed representations but omits the invariance term. Xiao et al. (2021) use contrastive learning with separate projection heads for each augmentation, treating same-augmentation pairs as negatives. ContextSSL (Gupta et al., 2024) conditions representations on both current actions and recent context using a memory module and employs a dual predictor for transformation prediction to avoid collapsing to invariance. Other approaches (Shakerinava et al., 2022; Gupta et al., 2023; Yerxa et al., 2024) do not require explicit transformation parameters but instead enforce equivariance by applying the same transformation to multiple view pairs and minimizing a distance-based loss. Finally, action-conditioned JEPAs like IWM incorporate augmentation parameters and mask positions by conditioning the predictor to induce equivariance without additional objectives, though they still face the invariance-equivariance tradeoff depending on predictor capacity (Garrido et al., 2024).
Positioning of Our Work. seq-JEPA belongs to the family of joint-embedding predictive architectures and is a non-generative world model. Unlike most equivariant SSL methods, seq-JEPA does not rely on an equivariance loss or transformation prediction objective, nor does it require view pairs with matched transformations. Instead, it introduces architectural inductive biases to JEPAs (Assran et al., 2023; Garrido et al., 2024) to learn separate invariant and equivariant representations, which emerge naturally from its action-conditioned sequential predictive learning. Furthermore, in contrast to ContextSSL, which extends the two-view contrastive setting of contrastive predictive coding van den Oord et al. (2019) using a transformer decoder projector conditioned on previous views via key-value caching, seq-JEPA operates on sequences of action-observation pairs in an online end-to-end manner by incorporating a sequence model (e.g. a transformer encoder) as a learned working memory during both training and inference. Moreover, while ContextSSL aims to adapt equivariance to recent transformations by dynamically modifying the training distribution, seq-JEPA is designed to explicitly learn both invariant and equivariant representations with respect to a specified set of transformations, and is also well-suited for downstream tasks that require multi-step observation aggregation (e.g., predictive learning across saccades and path integration).
6 Limitations and Future Perspectives
We have validated viability of seq-JEPA across a range of transformations in the image domain. Here, we discuss limitations and possible future directions. First, the transformer-based aggregator in seq-JEPA could support multi-modal fusion across language, audio, or proprioceptive inputs—enabling multi-modal world modeling and generalization. Second, while our results show clear benefits from longer sequences, we have experimented with relatively short training and inference sequence lengths as the ResNet backbone is also trained end-to-end. The sequence scaling trends observed in the paper suggest that seq-JEPA can benefit from longer context windows over representations of pre-trained backbones as in (Pang et al., 2023; Lin et al., 2024). Third, our method assumes access to a known transformation group (e.g., $SO(3)$ for 3D rotations). Designing group-agnostic or learned transformation models (Finzi et al., 2021) without access to transformation parameters or pairs of same transformation is an open challenge in equivariant SSL that future work may tackle. Finally, our preliminary ImageNet-1k transfer results (Appendix B.7) point to potential for broader generalization. Scaling seq-JEPA to larger foveated image settings or video and multi-modal datasets such as Song et al. (2023) could support the development of lightweight, saliency-driven agents capable of learning efficiently from partial observations in embodied settings with a limited field of vision.
Acknowledgments and Disclosure of Funding
This project was supported by funding from NSERC (Discovery Grants RGPIN-2022-05033 to E.B.M., and RGPIN-2023-03875 to S.B.), Canada CIFAR AI Chairs Program and Google to E.B.M., Canada Excellence Research Chairs (CERC) Program, Mila - Quebec AI Institute, Institute for Data Valorization (IVADO), CHU Sainte-Justine Research Centre, Fonds de Recherche du Québec–Santé (FRQS), and a Canada Foundation for Innovation John R. Evans Leaders Fund grant to E.B.M. This research was also supported in part by Digital Research Alliance of Canada (DRAC) and Calcul Québec.
Appendix A Implementation Details
A.1 Data preparation
3DIEBench. The original $256\times 256$ images are resized to a $128\times 128$ resolution for all experiments. Normalization is done using the means and standard deviations in Garrido et al. [2023], i.e., $\mu=[0.5016,0.5037,0.5060]$ and $\sigma=[0.1030,0.0999,0.0969]$ for the three RGB channels, respectively.
CIFAR100. We use $32\times 32$ images with normalization parameters typically used in the literature, i.e., $\mu=[0.4914,0.4822,0.4465]$ and $\sigma=[0.247,0.243,0.261]$. For data augmentation, we follow EquiMod’s augmentation strategy.
Tiny ImageNet. The training set consists of 100000 ImageNet-1k images from 200 classes (500 for each class) downsized to $64\times 64$. The validation set has 50 images per class. We use normalization parameters $\mu=[0.4914,0.4822,0.4465]$ and $\sigma=[0.247,0.243,0.261]$, for the three RGB channels, respectively. For data augmentation, we use the same augmentation parameters as CIFAR100 with the kernel size of Gaussian blur adapted to the $64\times 64$ images.
STL10. In order to extract the saliencies, we resize images to $512\times 512$, feed them to the pre-trained DeepGaze IIE [Linardos et al., 2021], resize the output saliencies back to $96\times 96$, and store them alongside original images. We use normalization parameters $\mu[0.4467,0.4398,0.4066]$, $\sigma=[0.2241,0.2215,0.2239]$ for the three RGB channels, respectively. After sampling fixations from saliencies, the patches that are extracted from the image to simulate foveation are $32\times 32$ (compared to the full image size of $96\times 96$). For IoR, we zero-out a circular area with radius of $16$ around each previous fixation.
Transformation parameters. For 3DIEBench, we use the rotation and color parameters provided with images for action conditioning as done in Garrido et al. [2023]. For the augmentation setting, we use the parameters corresponding to each of the three augmentations and form the action as the relative augmentation vector between two images. For crop, we use four variables, i.e., vertical and horizontal coordinate, and height and width. For color jitter, we use four variables: brightness, contrast, saturation, and hue. For blur, we use one variable: the standard deviation of the blurring kernel. In predictive learning across saccades, the action is a 2-d vector, i.e., the normalized relative $(x,y)$ coordinate between two patches.
A.2 Training and evaluation details
Additional Training Details.
We used the PyTorch framework for training all models. For experiments that use CIFAR100 and low-resolution STL-10 patches in predictive learning across saccades, we use the CIFAR variant of ResNet-18. For models trained with AdamW, we used with default $\beta_{1}$ and $\beta_{2}$, a weight decay of $0.001$, and a learning rate of $4\times 10^{-4}$ with a linear warmup for $20$ epochs starting from $10^{-5}$ followed by a cosine decay back to $10^{-5}$. For models trained with Adam, we use the Adam optimizer with a learning rate of $10^{-3}$, default $\beta_{1}$ and $\beta_{2}$, and no weight decay.
Protocols for training evaluation heads.
For linear probing, we follow a common SSL protocol and train a linear classifier on top of frozen representations with a batch size of 256 for 300 epochs using the Adam optimizer with default hyperparameters. For action prediction, we follow a similar protocol as SIE [Garrido et al., 2023]. Specifically, for rotation, color jitter, and crop, we train an MLP regressor with a hidden dimension of 1024 and ReLU activation for 300 epochs. For color (in 3DIEBench), blur (in the augmentation setting), and position (in predictive learning across saccades), we use a linear regressor and train it for 50 epochs. For path integration experiments, the same regressor architectures as the equivariance evaluation heads are used, i.e., the MLP for angular path integration and the linear regressor for saccade path integration. All regression heads are trained using the Adam optimizer with default hyperparameters.
Hardware. Each experiment was run on a single NVIDIA A100 GPU with 40GB of accelerator RAM. A single run of seq-JEPA on 3DIEBench with a training sequence length of three takes around 15 hours.
A.3 Architectural details
Below, we describe the architectural details and hyperparameters specific to each baseline.
BYOL.
We use a a projection head with 2048-2048-2048 intermediate dimensions. The predictor has a hidden dimension of 512-d with ReLU activation. We use the same EMA setup outlined in the original paper [Guo et al., 2020], i.e., the EMA parameter $\tau$ starts from $\tau_{\text{base}}=0.996$ and is increased following a cosine schedule.
SimCLR.
We use a temperature parameter of $\tau=0.5$ with a projection MLP with 2048-2048-2048 intermediate dimensions.
VICReg.
We use $\lambda_{\text{inv}}=\lambda_{V}=10$, $\lambda_{C}=1$, and a projection head with 2048-2048-2048 intermediate dimensions.
Conditional BYOL.
The architecture is the same as BYOL, except that the predictor also receives the normalized relative transformation parameters. We use a linear action projector of 128-d and the same EMA setup as BYOL.
SIE.
For both invariant and equivariant projection heads, we use intermediate dimensions of 1024-1024-1024. For the loss coefficients, we use $\lambda_{\text{inv}}=\lambda_{V}=10$, $\lambda_{\text{equi}}=4.5$, and $\lambda_{C}=1$. We use the hypernetwork architecture for all experiments.
SEN.
We use a temperature parameter of $\tau=0.5$ with a projection MLP with 2048-2048-2048 intermediate dimensions.
EquiMod.
We use the version based on SimCLR (both invariant and equivariant losses are contrastive with $\tau=0.1$ and have equal weights). The projection head has 1024-1024-128 intermediate dimensions. We use a linear action projector of 128-d.
ContextSSL.
We use the pre-trained weights provided by the authors (trained for 1000 epochs) and follow their evaluation protocol on 3DIEBench.
seq-JEPA.
For the sequence aggregator, we use a transformer encoder with three layers, four attention heads, and post-normalization. For the predictor, we use an MLP with a hidden layer of 1024-d and ReLU activation. The linear action projector in our default setting is 128-d. We use the same EMA setup as BYOL.
Appendix B Additional Experimental Results
B.1 Evaluation results on 3DIEBench for models conditioned on rotation and color
Table 5 reports performance of seq-JEPA and several equivariant baselines when conditioned on both rotation and color in 3DIEBench. All methods suffer a drop in classification performance and become highly sensitive to color. Similar performance degradations have been previously observed with 3DIEBench [Garrido et al., 2023, Gupta et al., 2024], though without a clear explanation. One plausible explanation aligned with Principle II in Wang et al. [2024], is that color in 3DIEBench, i.e., floor and light hue, are weakly correlated with class labels in 3DIEBench (low class relevance). Therefore, forcing the encoder to encode color information would cause class information to be lost, resulting in degradation of classification accuracy.
Table 5: Evaluation on 3DIEBench for rotation and color prediction (equivariance) and linear probe classification (invariance). All models are conditioned on both rotation and color.
Method Classification (top-1) Rotation pred. ($R^{2}$) Color pred. ($R^{2}$)
SEN 82.17 0.29 0.96
EquiMod 82.58 0.27 0.95
Conditional BYOL 81.95 0.38 0.94
SIE 75.34 0.46 0.97
seq-JEPA ($M_{tr}=4$, $M_{val}=5$) 79.31 0.52 0.97
B.2 Measuring invariance of aggregate representations
In order to have a more direct measure of invariance of the aggregate representation, we trained a model on 3DIEBench with $M_{val}=3$ and rotation conditioning, then attempted to predict the individual rotation parameters of each input view from $z_{\text{AGG}}$. To prevent leakage of rotation information during prediction, all action embeddings were zeroed out. The resulting mean $R^{2}$ score obtained by predicting individual rotations of the three views using $z_{AGG}$ is $0.141\pm 0.002$; indicating that the aggregate representation contains little to no rotation-specific information.
B.3 Complete evaluation results for linear probing on top of aggregate representations
We provide our complete evaluation results for linear probing on top of seq-JEPA’s aggregate representations for our three transformation settings with different training and inference sequence lengths. Figure 8 shows the top-1 accuracy on 3DIEBench models conditioned on rotation. Figure 9 shows top-1 accuracy on STL-10 for models trained via predictive learning across saccades. Figures 10 and 11 show top-1 accuracy on CIFAR100 and Tiny ImageNet, respectively, with different types of action conditioning (crop, color jitter, blur, or all three). These heatmaps reflect the same trends observed in Figure 7 and discussed in Section 4.6, illustrating the consistent effect of sequence length on representation quality.
Figure 8: seq-JEPA’s performance on 3DIEBench with rotation conditioning; the heatmap shows linear probe accuracy on top of aggregate representations for different training and inference sequence lengths.
Figure 9: seq-JEPA’s performance on STL-10 with predictive learning across saccades; the heatmap shows linear probe accuracy on top of aggregate representations for different training and inference sequence lengths.
Figure 10: seq-JEPA’s performance on CIFAR100 with different types of action conditioning (crop, color jitter, blur, or all three); the heatmap shows linear probe accuracy on top of aggregate representations for different training and inference sequence lengths.
Figure 11: seq-JEPA’s performance on Tiny ImageNet with different types of action conditioning (crop, color jitter, blur, or all three); the heatmap shows linear probe accuracy on top of aggregate representations for different training and inference sequence lengths.
B.4 Comparison of evaluation results on encoder representations and aggregate representations
For completeness, we provide linear probe classification on encoder representations for different transformation settings in Table 6 and compare them with accuracy on aggregate representations for different inference evaluation lengths. The aggregate representation generally achieves a much higher classification performance thanks to the architectural inductive bias in seq-JEPA.
Table 6: Comparison of seq-JEPA classification performance across datasets and conditioning. Top-1 classification accuracy is reported for $z_{res}$ and $z_{agg}$, with varying inference lengths $M_{eval}$.
Dataset Conditioning $M_{tr}$ $z_{res}$ $z_{agg}$
$M_{eval}=1$ $M_{eval}=3$ $M_{eval}=5$
3DIEBench None 3 80.91 81.61 86.05 87.36
3DIEBench Rotation 1 84.88 84.08 85.34 85.31
3DIEBench Rotation 3 82.49 81.72 85.32 87.41
3DIEBench Rotation + Color 4 74.88 71.14 75.97 79.31
CIFAR100 None 2 53.00 51.60 57.05 58.37
CIFAR100 Crop + Jitter + Blur 1 56.23 54.24 52.90 52.92
CIFAR100 Crop + Jitter + Blur 2 52.07 53.07 59.34 60.35
CIFAR100 Crop + Jitter + Blur 3 46.31 49.48 57.60 58.33
CIFAR100 Crop 2 52.62 52.06 58.07 59.32
CIFAR100 Color Jitter 3 54.92 50.21 57.20 58.62
CIFAR100 Blur 3 51.41 48.69 55.54 56.72
Tiny ImageNet None 2 32.84 30.48 35.03 35.97
Tiny ImageNet Crop + Jitter + Blur 1 33.03 32.34 36.14 37.07
Tiny ImageNet Crop + Jitter + Blur 2 27.20 30.57 34.99 35.56
Tiny ImageNet Crop + Jitter + Blur 3 24.74 28.84 33.78 34.68
Tiny ImageNet Crop 2 31.13 30.89 35.07 35.69
Tiny ImageNet Color Jitter 3 31.85 29.05 34.27 35.21
Tiny ImageNet Blur 3 27.20 28.83 34.82 35.79
STL-10 None 4 61.38 62.21 69.06 70.45
STL-10 Position 4 81.20 71.45 81.53 83.44
STL-10 Position (no saliency) 4 79.29 63.14 76.93 79.85
STL-10 Position (no IoR) 4 72.49 68.95 76.84 77.97
B.5 Additional UMAP Visualizations
In Figure 12, we visualize the UMAP projections of seq-JEPA representations trained on 3DIEBench without action conditioning. Similarly, Figure 14 shows projections for models trained on STL-10 via predictive learning across saccades, also without action conditioning. Compared to the action-conditioned counterparts (Figures 5 and 13), these projections exhibit weaker or no smooth color gradients in their corresponding transformation-colored UMAP—indicating reduced equivariance to transformation parameters.
Figure 12: 2-D UMAP projections of seq-JEPA representations on 3DIEBench without action conditioning ($M_{tr}=3$, $M_{val}=5$). Encoder outputs are color-coded by class (left) and rotation angle (middle); aggregate token representations are color-coded by class (right).
Figure 13: 2-D UMAP projections of seq-JEPA representations on STL-10 with action conditioning ($M_{tr}=M_{val}=4$). Encoder outputs are color-coded by class (top-left) and by $X$/$Y$ fixation coordinates (bottom); the aggregate token is color-coded by class (top-right).
Figure 14: 2-D UMAP projections of seq-JEPA representations on STL-10 without action conditioning ($M_{tr}=M_{val}=4$). Encoder outputs are color-coded by class (left) and fixation coordinates (middle); the aggregate token is color-coded by class (right).
B.6 Details of Path Integration Experiments.
While an agent executes a sequence of actions in an environment, transitioning from an initial state to a final state, it should be capable of tracking its position by integrating its own actions. This is also a crucial cognitive ability that enables animals to estimate their current state in their habitat [McNaughton et al., 2006]. Here, we evaluate whether seq-JEPA is capable of path integration. Given the sequence of observations $\{x_{i}\}_{i=1}^{M+1}$ generated from transformations $\{t_{i}\}_{i=1}^{M+1}$ and the corresponding relative action embeddings $\{a_{i}\}_{i=1}^{M}$, we define the task of path integration over the sequence of actions as predicting the relative action that would directly transform $x_{1}$ to $x_{M+1}$ given $z_{AGG}$ and $a_{M}$. In other words, given the aggregate representation of a sequence of action-observation pairs and the next action, we would like to predict the overall position change from the starting point ($x_{1}$) to the end point ($x_{M+1}$). We consider path integration for rotation angles in 3DIEBench and across eye movements with STL-10. For rotations, the task is integrating a series of object rotations from the first view to the last, i.e. angular path integration. For eye movements, the task is integrating the eye movements from the first saccade to the last, i.e. visual path integration. To measure path integration performance for inference sequence length $M$, we train a regression head on top of the concatenation of $z_{AGG}$ and $a_{M}$ to predict the transformation from $x_{1}$ to $x_{M+1}$. Figure 6 shows that seq-JEPA performs well in both angular and visual path integration. The red curve corresponds to the performance of the original seq-JEPA. The blue curve corresponds to experiments in which the action embeddings are ablated (zeroed-out during inference for all views). The green curve corresponds to experiments in which the encoder (visual) representations are ablated during inference. As expected, path integration becomes more difficult as the number of observations increases (red curves). Ablating action conditioning (blue curves) results in failure of path integration. On the other hand, ablating the visual representations (green curves) results only in a small performance drop compared to the original model, indicating that action conditioning is the key factor that enables path integration.
B.7 Transfer learning results on ImageNet-1k
To evaluate generalization of our model beyond STL-10, we assess transfer performance of the model trained on STL-10 patches via predictive learning across saccades on ImageNet-1k. We follow the same linear probing protocol as in-distribution evaluations: we freeze the ResNet and transformer encoder and train a linear classifier on aggregate representations generated from foveated patches.
We extract saliency maps for ImageNet-1k images using DeepGaze IIE, resize them to $224\times 224$, and sample sequences of patches sized $32\times 32$ or $84\times 84$ for this out-of-distribution (OOD) evaluation setting. Figure 15 shows top-1 linear probe accuracy on ImageNet-1k validation set across varying inference sequence lengths ($M_{val}$). For both patch sizes, performance improves with longer inference sequences, validating seq-JEPA’s ability to benefit from extended context even in this difficult OOD ImageNet-1k setting. These results echo findings in the main experiments for different training and inference sequence lengths and confirming the possibility of model’s scalability in terms of data, parameter count, and compute.
Figure 15: Linear probe transfer learning accuracy on ImageNet-1k for two different patch sizes; the model is pre-trained on STL-10 via predictive learning across saccades.